(need help with problem e.)

Terri Vogel, an amateur motorcycle racer, averages 129.71 seconds per 2.5 mile lap (in a seven-lap race) with a standard deviation of 2.28 seconds. The distribution of her race times is normally distributed. We are interested in one of her randomly selected laps.
c. Find the percent of her laps that are completed in less than 130 seconds (write answer as a decimal). Round answer to three decimal places (i.e. 0.123) Answer

0.551


d. The fastest 3% of her laps are under how many seconds? Round answers to one decimal place (i.e. 12.3) Answer

125.4


e. The middle 80% of her laps are from Answer

__126.8___ seconds to _______ seconds?

Round answers to one decimal place (i.e. 132.6).
Answer #1

c) The  percent of her laps that are completed in less than 130 seconds is

130-129.71 PZ 0.13) 0.551-55.1% from area under the standard normal table d) X=mean-23%SD = 129.71-1.88 1 * 2.28 = 125.4 nce Z3%-1881 from area under the standard normal table e) From area under the standard normal curve (or Z-Tables,) P(-1.282 <Z < 1.282) = 0.80 (1) Now,we can find the a and b such that middle 80% of her laps P(a X<b)-0.80 a-129.717 b-12971-0.80.(2) <z < 2.28 2.28 From() and (2) a-129.71 -1.282 |a-126.81 2.28 b-19.11.2821326 2.28